OpenStudy (anonymous):

Perform indicated operation and, is possible, simplify. If quotient is undefined state this. Multiply in the denominator before dividing by that value. (-2)(5)/(7)(-8)

OpenStudy (agreene):

\[\frac{-2*5}{7*-8}=\frac{-10}{-56}=\frac{10}{56}=\frac{5}{28}\approx0.17857142\] the last 6 digits of that number are repeating.